OpenStudy (anonymous):

determine the inputs that yield the minimum value: F(x)=sqrt[3]{2x^2-8x+22}?????

myininaya (myininaya):

have you had calculus?

OpenStudy (anonymous):

no..i am in precal

myininaya (myininaya):

\[f(x)=\sqrt{3}*(2x^2-8x+22)\] \[f'(x)=\sqrt{3}(4x-8)\] set f' equal to zero \[\sqrt{3}(4x-8)=0, x=2\]

myininaya (myininaya):

ok so since you haven't cal i will do this algebra way also

OpenStudy (anonymous):

\[F(x)\sqrt[3]{2x^2-8x+22} \] just wanna make sure we r both seeing the same problem

myininaya (myininaya):

\[f(x)=\sqrt{3}*2(x^2-4x)+\sqrt{3}*22=2\sqrt{3}(x^2-4x+(\frac{-4}{2})^2)+22\sqrt{3}-2\sqrt{3}*(\frac{-4}{2})^2\] \[=2\sqrt{3}(x^2-4x+4)+(22-2)\sqrt{3}=2\sqrt{3}(x-2)^2+20\sqrt{3}\] so we get the min x=2

myininaya (myininaya):

so you have another probem to \[f(x)=\sqrt[3]{2x^2-8x+22}\]

myininaya (myininaya):

so you don't know how to do the derivative right?

OpenStudy (anonymous):

i do not

OpenStudy (anonymous):

this was the problem from the beginning

OpenStudy (anonymous):

so is the min. value h?

myininaya (myininaya):

\[f(x)=\sqrt{3}(2x^2-8x+22)\]

OpenStudy (anonymous):

minimum value will be at the vertex of the quadratic. forget the cube root. the quadratic is a parabola that faces up. will have a minimum value at \[x=-\frac{b}{2a}\]

myininaya (myininaya):

x=2 is the min then

OpenStudy (anonymous):

you have \[2x^2-8x+22\] \[-\frac{b}{2a}=\frac{8}{4}=2\]

OpenStudy (anonymous):

no my myininaya 2 is not the minimum value

OpenStudy (anonymous):

2 is the input that will get the minimum value yes? minimum means minimum of function, not x that gives it

OpenStudy (anonymous):

yessss!! i get as far as getting the 2 but when i plug in i must do something wrong bcuz my answer is the matching up with the correct answer

OpenStudy (anonymous):

\[F(2)=\sqrt[2]{2\times 2^2-8\times 2+22}\] \]

OpenStudy (anonymous):

rather the cube root. i get the cube root of \[8-16+22=-8+22=14\] so minimum should be \[\sqrt[3]{14}\]

OpenStudy (anonymous):

2.41 rounded

OpenStudy (anonymous):

oh lordamercy i guess i should read the problem. it says determine the inputs that yield the minimum value:

OpenStudy (anonymous):

it wants the input. and the input is 2!

OpenStudy (anonymous):

the book says (2,2.4) so u were still right.

OpenStudy (anonymous):

jeez louise

myininaya (myininaya):

i didn't say the minumum value was 2 i said the minumum was 2

OpenStudy (anonymous):

?

myininaya (myininaya):

the minumum value is 20sqrt{3}

OpenStudy (anonymous):

achieves a minimum at (2,2.4)

OpenStudy (anonymous):

my master, minimum value is the same as the minimum. output, not input. on the other hand problem asked for input, so i guess it wanted 2 as the answer.

OpenStudy (anonymous):

if that is what the book says the book is trash

OpenStudy (anonymous):

minimum value means just that. the minimum value. not a coordinate

myininaya (myininaya):

no the value refers you the y

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

cant argue with u there satel, ive found a hand full of mistakes in this book, so the answer would b 2 afterall?

myininaya (myininaya):

the min value would be 20sqrt{3}

OpenStudy (anonymous):

the answer to the question "determine the inputs that yield the minimum value?" is 2

OpenStudy (anonymous):

the answer to the question "what is the minimum value?" is \[\sqrt[3]{14}\]

myininaya (myininaya):

i honeslt dont know what function we are looking at there has been mention of two different functions \[f(x)=\sqrt{3}(2x^2-8x+22)\] and \[f(x)=\sqrt[3]{2x^2-8x+22}\]

OpenStudy (anonymous):

the answer to "what is the coordinate of the lowest point on the curve of \[y=F(x)\] ?" is\[ (2,2.41)\]

myininaya (myininaya):

i did the first one

OpenStudy (anonymous):

@mininaya i have been assuming it is the second one

OpenStudy (anonymous):

no myin it was the same function we didnt understand e/o at the beg.

myininaya (myininaya):

so it was the first one

OpenStudy (anonymous):

it is the 2nd

OpenStudy (anonymous):

\[F(x)=\sqrt[3]{2x^2-8x+22}\] yes?

OpenStudy (anonymous):

ha ha satael73 u were doin the correct one, myi we had a misunders.

OpenStudy (anonymous):

the input would have been the same in either case for the same reason. the output would of course be different

OpenStudy (anonymous):

either way sate. i believe my professor would like the answer (2,2.4) so i thank u 4 wrkin it out step by step

myininaya (myininaya):

i wish i knew what problem we were doing lol

OpenStudy (anonymous):

ha ha myin. i thank u too! i appreciate it!

OpenStudy (anonymous):

you are welcome, so long as you tell your professor that you have it on good authority that "minimum" or "maximum" of a function refers to an output, not a coordinate!

OpenStudy (anonymous):

my mininaya please tell me you haven't been here since 8 this morning!

OpenStudy (anonymous):

ha ha i will do so, just as soon as my semester is over and my grade is finalized ha ha

myininaya (myininaya):

i always understood if you don't say the word value you are talking about the x if you talking about value , than you are talking about y